OpenStudy (anonymous):

Solve the system of equations. 5x+1y=27 x=-2y

OpenStudy (anonymous):

subtitute value of x=-2y in first equation 5(-2y) +1y =27 -10y +y = 27 -9y =27 ((y = -3)) substitute this value of y in second equation. x = -2 (-3) ((x = 6)) {(6,-3)}